Physical Chemistry

30gm of an aqueous solution of a particle solute (containing 30% by mass solute ) is mixed with 400gm of another aqueous solution of same solute ( containing 40% solute by mass ). In the final solution calculate .
( Given: density of final solution= 7/8 gm/ ml, Molecular mass of solute= 50)
  1. mass% of solute
  2. mole fraction of solute
  3. molarity
  4. mmolality

AMIT RELAN
11 Years agoGrade

1 Answer

Naveen Kumar
11 Years ago
Consider about first solution,
mass of solution=30g
massof solute= 30% of 30g=30*30/100=9g
consider another solution. In this case
mass of solution=400g
mass of solute=400*40/100=160g
Now mixing both the solution,
mass of solution=400+30g=430g
mass of solute=160+9 g=169g
Now, volume of solution=mass/density=430/(7/8) ml
now calculate the quatities asked.
% of mass of solute in final solution=(169/430)*100
now calculate the moles of water and of solute by the formula,
Moles=mass/molecular-mass
mole of water=m(say)and moles of slute=n(say)
so mole fraction=n/(m+n)
to calculate molarity ,
molarity=n/V *1000
molality=n/18*m n*1000
to know how to calculate these terms and how to use these concepts in chemistry, please go through a good text-book .
If you are getting any doubt, u can post on nforum.